OpenStudy (anonymous):

seperate the seprebale equation (ds/dt) = (4.25)-(3s/200)

OpenStudy (anonymous):

Calc II question

OpenStudy (anonymous):

Someone help please!

OpenStudy (anonymous):

you mean how to find ds/dt right ?

OpenStudy (anonymous):

no sir i do not want a derrivative

OpenStudy (anonymous):

\[\int\limits_{}^{}ds/(4.25-3s/200)=\int\limits_{}^{}dt\] I believe that separates it. The answer will be an exponential with e. Hope that helps some.

OpenStudy (anonymous):

(ds/dt) = (4.25)-(3s/200) take lcm on right hand side =>(ds/dt)=(480-3s)/200 =>ds/(480-3s)=dt/200 now integrate it on both sides =>[log(480-3s)]/-3=t/200

OpenStudy (anonymous):

GOT IT